Disc 2, Bb# 3 In the morning glory plant, three genes are linked as follows: 20 m.u 30 m.u A morning glory plant with the genotype BB gg DD was crossed to a morning glory plant of genotype bb GG dd. The F1 progeny was testcrossed to a b g d/bg d plant. A. Assume no interference, what genotypic classes will be found in the progeny and the frequencies if there are 1000 progeny? B. Assume 20% interference, what genotypic classes will be found in the progeny and the frequencies if there are 1000 progeny?

please show all work clearly
Answer #1

Answer:

BgD/BgD x bGd/bGd----Parents

BgD/bGd ------------------F1

BgD/bGd x bgd/bgd

A).

Expected double crossover progeny = 20% * 30% = 0.06

Class of gametes

Phenotype

Frequency of reciprocal pair

Numbers

Frequency of each class

Number of progeny frequency * 1000 = %

Parental

BgD

1- all recombinant progeny

1-(0.14+0.24+0.06) = 0.56

0.28

280 (BgD/bgd)

bGd

0.28

280 (bGd/bgd)

SCO 1

BGd

RF in region 1 = SCO1 – DCO

SCO1 = 0.2 – 0.06 = 0.14

0.07

70 (BGd/bgd)

bgD

0.07

70 (bgD / bgd)

SCO 2

Bgd

RF in region 2 = SCO2 – DCO

SCO2 =0.3 – 0.06 = 0.24

0.12

120 (Bgd/bgd)

bGD

0.12

120 (bGD / bgd)

DCO

BGD

(RF in region 1) * (RF in region 2)

0.06

0.03

30 (BGD/bgd)

Bgd

0.03

30 (bgd/bgd)

B).

Interference = 1- COC

20% or 0.2 = 1 – COC

COC = 1 – 0.2 = 0.8

COC = Observed double crossover progeny / Expected double crossover progeny

Expected double crossover progeny = 20% * 30% = 0.06

0.8 = Observed double crossover progeny / 0.06

Observed double crossover progeny = 0.06 * 0.8 = 0.048

Class of gametes

Phenotype

Frequency of reciprocal pair

Numbers

Frequency of each class

Number of progeny frequency * 1000 = %

Parental

BgD

1- all recombinant progeny

1-(0.152+0.252+0.048) = 0.548

0.274

274 (BgD/bgd)

bGd

0.274

274 (bGd/bgd)

SCO 1

BGd

RF in region 1 = SCO1 – DCO

SCO1 = 0.2 – 0.048 = 0.152

0.076

76 (BGd/bgd)

bgD

0.076

76 (bgD / bgd)

SCO 2

Bgd

RF in region 2 = SCO2 – DCO

SCO2 =0.3 – 0.048 = 0.252

0.126

126 (Bgd/bgd)

bGD

0.126

126 (bGD / bgd)

DCO

BGD

(RF in region 1) * (RF in region 2)

0.048

0.024

24 (BGD/bgd)

Bgd

0.024

24 (bgd/bgd)